# Learning with pyCub: A New Simulation and Exercise Framework for Humanoid Robotics

### v1

@@ Abstract
We present pyCub, an open-source physics-based simulation of the humanoid robot iCub, along with exercises to teach students the basics of humanoid robotics. Compared to existing iCub similators (iCub SIM, iCub Gazebo), which require C++ code and YARP as middleware, pyCub works without YARP and with Python code. The complete robot with all articulations has been simulated, with two cameras in the eyes and the unique sensitive skin of the iCub comprising 4000 receptors on its body surface. The exercises range from basic control of the robot in velocity, joint, and Cartesian space to more complex tasks like gazing, grasping, or reactive control. The whole framework is written and controlled with Python, thus allowing to be used even by people with small or almost no programming practice. The exercises can be scaled to different difficulty levels. We tested the framework in two runs of a course on humanoid robotics. The simulation, exercises, documentation, Docker images, and example videos are publicly available at https://rustlluk.github.io/pyCub.
I Introduction
@@
There is a number of good robotics textbooks, such as [1], or [2] with a great video supplement, but few of them offer hands-on exercises, toolboxes, or robot simulators. The book of Corell et al. [3] is accompanied with a github repository with Matlab and Mathematica code. Peter Corke developed a Python Robotics Toolbox [4]. This toolbox provides tools for both dynamics and kinematics of robots and is a successor of Robotic Toolbox written in MATLAB [5]. Tutorials on robot differential kinematics using the Python Robotics Toolbox were presented by Haviland and Corke [6, 7]. Sakai et al. [8] proposed an educational textbook with Python examples to explain several robotic algorithms for autonomous navigation. One part of our work is similar in the sense that we provide exercises that should teach core robotics.
@@
Figure 1: An example from the simulation environment showing the iCub humanoid robot with artificial skin (red dots), a table and a red ball that can be manipulated by the robot.
@@
There are other toolboxes for kinematics and dynamics [9, 10] or self-calibration of robots [10, 11]. As many robots are controlled and unified under Robot Operating System (ROS), there are frameworks based on this middleware. The authors in [12] created a ROS toolbox for simulating robots with compliant actuators. Fischer et al. [13] proposed a framework that combines Conda and Jupyter Notebooks for a more accessible introduction to ROS. Another interesting project is Unibotics [14] that provides web-based learning platform mainly for navigation and mapping. Furthermore, there are more specialized toolboxes, e.g., for modular industrial robots [15] or aerial robotic manipulators [16].
@@
Our course is on humanoid robots, but due to the universal nature of such a platform, it can be used to teach several robotics topics, from kinematics and dynamics, over walking, gaze control, grasping, to human-robot interaction. Our objective thus was a robot simulator easily accessible by students that can be used for hands-on exercises from the course. We selected iCub [17] as it is a full humanoid and an open-source research platform. The dimensions of the robot resemble a four-year-old child and the geometry and kinematics is designed to mimic the human body as much as possible (at the joint level, not at the muscle level). There are 53 active Degree of Freedoms. A unique feature of iCub is the whole-body artificial skin with more than 4000 individual tactile sensors. The robot has a legacy simulator “iCub SIM” [18] and a newer simulator in Gazebo [19] and offers its own training course [20]. However, getting the environment running is an intricate process, and the default control is in C++ using a Yet Another Robot Platform (YARP[21] middleware, which is not suitable for beginners. After having run two editions of the course on the iCub Gazebo, with Docker containers and other infrastructure prepared, WE have evaluated the feedback forms and realized that the students complained about the complexity of the whole system. Thus, we decided to simplify the whole pipeline and create a simulator using Python, which is a more user-friendly language. The comparison of existing iCub simulators is in Table I. We are not aiming to replace the existing simulators, but to provide a more accessible way to iCub and humanoid robotics in general.
iCub SIM iCub Gazebo pyCub
@@
Python
Middleware YARP YARP None
Transferable
to real iCub
Yes Yes No
Skin Yes Yes* Yes
High-level
planners
Yes** Yes** No
Physics
Engine
ODE DART*** Bullet
Rendering
Engine
OpenGL OGRE
@@
TABLE I: Comparison of pyCub with existing iCub simulators (iCub SIM [18] and iCub Gazebo [19]). *Skin to Gazebo was added later by the community. **The planners are coming from the connection with YARP. ***By default, can be changed.
We present pyCub, a Python physics-based simulator of the humanoid robot iCub. The simulator contains all limbs of the robot, together with the possibility to visualize view from the robot’s cameras and also with simulated artificial whole-body skin consisting of more than 4000 individual tactile sensors. The framework provides easy-to-use functions to control the robot in joint, velocity, and Cartesian space, and effortless access to proprioceptive, visual, and tactile sensors of the robot. Everything is written and accessible in Python, allowing even less advanced users to operate. The environment is fully customizable and allows to insert other objects—see Fig. 1 for an example. In addition to the simulator, we provide exercises of different difficulty levels ranging from basic control and movements to tasks typical for humanoid robots such as gazing, grasping, or whole-body skin processing and obstacle avoidance. The exercises can be further customized and allow difficulty adjustments and thus are suitable for learning (humanoid) robotics on different difficulty levels. The framework was deployed in two runs of a course on humanoid robotics.
@@ II Simulation Environment
II-A Technical Details
The physics-backbone of the simulator is realized with the PyBullet engine [22]. This module was selected because it is open-source, multi-platform and does not require excessive computational resources—mainly compared to recent physical simulators such as Orbit (NVIDIA Issac Sim) [23]. The average performance with Graphical User Interface (GUI) and skin enabled is around real time [^1] Tested on Ryzen 7 PRO 4750u CPU.. The most computationally demanding aspects are the number of DoFs and the model used, which is very complicated and precise. Thus, when using different models, the performance can be much better. We provide a simplified model (without fingers; simpler meshes; etc.), with which the performance is boosted two times—while the model can be further simplified. For example, we created a simulation environment based on pyCub for a different project where a 6-DoF manipulator is used and the performance is in orders of magnitude of real time, which is suitable even for reinforcement learning algorithms.
The schematic of the simulator package is shown in Fig. 2.
@@ II-A Technical Details
II-A1 Time Management
The simulation is not running by itself, but the users have to perform every step manually, i.e., call update_simulation(). It is advantageous because even when using less powerful computers there is enough time to check the state of the world and perform computations between simulation steps. Additionally, there is an option to actually perform the simulation step only if a predefined duration elapsed from the last one. It is useful to slow down the time to, for example, check movements in more detail. The GUI has its own time management, as updating it with every simulation would make everything slower without benefits, as one simulation step does not visibly change the world state.
II-A2 GUI
@@
Figure 3: Example from the native GUI with output from both eyes and addition visualization showing vectors where the robot is looking (blue) and where it should be looking (red) used in the gaze exercise.
II-A3 Customizations
The environment is fully customizable through .yaml configuration files. The files are loaded when initializing an instance of the simulator, and thus users can create different configurations for different exercises/tasks/demos. The basic options allow to enable/disable GUI, artificial skin, eye-cameras output, logging, or self collisions. Further, the configurations allow us to set an initial position for every joint of the robot, the end-effector used for Cartesian control, or simulation step size. Finally, users can insert additional objects into the simulation. The configs support Unified Robot Description Format (URDF) files, or directly .obj triangular-mesh files. One can further specify the initial position, color, and whether the object is subject to gravity and can be moved (e.g, ball) or not (e.g., table).
II-A4 Installation
To provide easy installation, the framework can be installed from PyPI. Alternatively, we provide several versions of prebuilt Docker images or version runnable entirely online using the GitPod [^2] https://www.gitpod.io/ platform. Everything is available at https://rustlluk.github.io/pyCub.
II-B Control Modes
@@
[…]
@@
II-C Artificial Skin
The artificial skin is an integral component of the robot. However, the real robot contains more than 4000 individual sensors, which makes the simulation challenging. The sensors are also not binary, i.e., they can return value with 8-bit resolution. We decided to simulate the skin using raycasting. From every skin sensor (taxel) we cast a short ray with a given maximum length. If the ray hits something (other part of the robot or object in the environment), the output of the sensor is computed proportionally to the length of the ray. Using this method, we are able to simulate the intensity of contact and also the fact that for the real skin, even taxels around the touch point are lightly activated because of the deformation of the skin. The raycasting needs to be done in every step of the simulation, and, even though the raycasting is optimized and done in batch, it is still computationally demanding. Thus, we decided to first check (for every skin part) whether there is a close object (using overlaps of bounding boxes of objects) and perform the raycasting only if necessary. The values for each activated taxels are saved in a dictionary inside the simulator and binary activations are shown as green points in the visualization—see Fig. 4.
@@ II-C Artificial Skin
III Exercises
This Section introduces and describes the exercises available in the framework. We prepared several exercises that, in our opinion, span important tasks in robotics. We will describe those in recommended order based on their difficulty. However, the exercises can be further adjusted to change the difficulty level. The exercises always contain a configuration file, assignment script and testing script. Videos of each exercise are available at https://rustlluk.github.io/pyCub and in the accompanying video.
III-A Push the Ball!
The first exercise is purely made with the purpose of learning how to control the simulator. The goal is to get a ball as far from a table as possible—see Fig. 1 for visualization. Basically, every imaginable approach is allowed. Ranging from hitting, pushing, or grasping and throwing the ball. Upon completion, users should be able to control the robot in some of the modes and understand basic functionalities. During this exercise, the self-collisions are turned off—it is visible in some examples in the accompanying video.
The testing script for this exercise computes distance of a ball and a table after two seconds after the last movement of the robot and provides distance in meters.
III-B Smooth Movements
As described in Section II-B, the Cartesian interface of the simulator does not include any underlying planner. This exercise aims to provide insight on how to move in a straight line or how to “draw” a shape, e.g., a circle. In addition, the movements should be “smooth”, i.e., a continuous movement without any jerks.
@@ III-B Smooth Movements
The testing script computes distance of the performed trajectory to a reference one. There are different metrics for line and circle movements. However, the smoothness of the movements must be checked manually.
III-C Gaze
Gaze control is a key ability such that the robot can bring some objects into the field of view. The goal of gaze can be described as focusing on the object of interest using, mainly, joints in the head and eyes. There is a gaze control implemented in the real iCub robot [25] that is capable of fixing to a given 3D using all joints in the neck and eyes. In this exercise, the problem is simplified to fixating the view of the robot using neck joints to a ball moving in XY-plane.
@@ III-C Gaze
The described solution is only one of many and it can be solved in both simpler and more complicated ways. Also, the thresholds for allowed error can be adjusted, which can make some simpler solutions infeasible.
@@
III-D Reactive Control and Skin Processing
The fourth exercise is focused on reactive control in the case of contact with the environment. It is an important aspect of Human-Robot Interaction (HRI), since ideal humanoid robots should be able to operate in close contact with humans, which inevitably leads to collisions.
@@ III-D Reactive Control and Skin Processing
where $\mathbf{J(q)}^{*}$ can be either transposition $\mathbf{J(q)}^{T}$, inverse $\mathbf{J(q)}^{-1}$, or Moore-Penrose pseudoinverse $\mathbf{J(q)}^{+}$. One of the educational points of the exercise is to test the options, as each of them has its pros and cons. The classic inverse can only be used when the Jacobian is square and has full rank. Pseudoinverse can be unstable near singular configurations and move the end-effector in a straight line. Transpose is easier to compute, but does not necessarily move in a straight line. In the exercise, some limbs are in a configuration where not all of the options provide the desired movement.
The exercise has different levels of difficulty depending on the limb in contact and the number of simultaneous contacts.
The testing script here runs the different levels of difficulty. At the moment, there is no automatic evaluation and performance must be checked manually.
An example of a more complex reactive controller based on differential kinematics for the real iCub robot was presented in [27].
III-E Grasping
This exercise is simplified by the fact that we always want to grasp the same green ball—the only thing that changes is the position of the ball on a table. The aim is also to practice core computer vision, and thus the users are guided to find the ball in the image plane (in an RGB image from one of the eyes) and obtain 2D center of the ball. The center can be deprojected to 3D given the center (u, v pixels in an image), the depth value for the tuple of pixels, and camera parameters (width; height; focal length). Then, the robot should move above the ball, get fingers to correct start position, move toward the ball, grasp it, and move it up to prove a stable grasp. An example of successful grasp can be seen in Fig. 5.
Figure 5: Successful grasp from the grasping exercise with view from the left eye.

### v2

# Learning with pyCub: A Simulation and Exercise Framework for Humanoid Robotics

@@ Abstract
We present pyCub, an open-source physics-based simulation of the humanoid robot iCub, along with exercises to teach students the basics of humanoid robotics. Compared to existing iCub simulators (iCub SIM, iCub Gazebo), which require C++ code and YARP as middleware, pyCub works without YARP and with Python code. The complete robot with all articulations has been simulated, with two cameras in the eyes and the unique sensitive skin of the iCub comprising 4000 receptors on its body surface. The exercises range from basic control of the robot in velocity, joint, and Cartesian space to more complex tasks like gazing, grasping, or reactive control. The whole framework is written and controlled with Python, thus allowing to be used even by people with small or almost no programming practice. The exercises can be scaled to different difficulty levels. We tested the framework in two runs of a course on humanoid robotics. The simulation, exercises, documentation, Docker images, and example videos are publicly available at https://rustlluk.github.io/pyCub.
keywords:
humanoids, robotics, robotics learning
1 Introduction
@@
There are a number of good robotics textbooks, such as [25] or [13] with a great video supplement, but few of them offer hands-on exercises, toolboxes, or robot simulators. The book of Corell et al. [3] is accompanied by a GitHub repository with Matlab and Mathematica code. Peter Corke developed a Python Robotics Toolbox [2]. This toolbox provides tools for both dynamics and kinematics of robots and is a successor to the Robotic Toolbox written in MATLAB [1]. Tutorials on robot differential kinematics using the Python Robotics Toolbox were presented by Haviland and Corke [10, 9]. et al.Sakai [24] proposed an educational textbook with Python examples to explain several robotic algorithms for autonomous navigation. One part of our work is similar in the sense that we provide exercises that should teach core robotics.
@@
Figure 1: An example of the simulation environment showing the iCub humanoid robot, a table, a green ball that can be manipulated by the robot, and a robotic skin (red points in the left image). The output from the left eye (in the right image) shows a successful grasp.
@@
There are other toolboxes for kinematics and dynamics [12, 19] or self-calibration of robots [19, 23]. As many robots are controlled and unified under Robot Operating System (ROS), there are frameworks based on this middleware. The authors in [14] created a ROS toolbox for simulating robots with compliant actuators. et al.Fischer [7] proposed a framework that combines Conda and Jupyter Notebooks for a more accessible introduction to ROS. Another interesting project is Unibotics [21], which provides a web-based learning platform mainly for navigation and mapping. Furthermore, there are more specialized toolboxes, e.g., for modular industrial robots [11] or aerial robotic manipulators [27]. Recently, a framework containing simulation of multiple manipulators, mobile and humanoid robots was proposed to unify robot learning [8].
@@
Our presented framework was created as a tool for our university course. Our course is on humanoid robots, but due to the universal nature of such a platform, it can be used to teach several robotic topics, from kinematics and dynamics, to walking, gaze control, grasping, and human-robot interaction. Our objective thus was to create a robot simulator that is easily accessible to students and can be used for hands-on exercises of the course. We selected iCub [16] as it is a full humanoid and an open-source research platform. The dimensions of the robot resemble a four-year-old child, and the geometry and kinematics are designed to mimic the human body as much as possible (at the joint level, not at the muscle level). There are 53 active Degree of Freedoms. A unique feature of iCub is the whole-body artificial skin with more than 4000 individual tactile sensors. The robot has a legacy simulator “iCub SIM” [26], a newer simulator in Gazebo [17] and offers its own training course [20]. However, getting the environment running is an intricate process, and the default control is in C++ using a Yet Another Robot Platform (YARP[15] middleware, which is not suitable for beginners—there are optional Python bindings available, but as these are created from the C++ code, the syntax can be hard to grasp for less experienced users. After running two editions of the course using the iCub Gazebo, even with Docker containers and other infrastructure prepared, we have evaluated the feedback forms and realized that the students complained about the complexity of the whole system. Thus, we decided to simplify the whole pipeline and create a simulator using Python, which is a more user-friendly language. The comparison of existing iCub simulators is in Table 1—the original simulators and a neuromorphic (and event-based) simulator [6]. We are not aiming to replace the existing simulators, but to provide a more accessible way to iCub and humanoid robotics in general.
Table 1: Comparison of pyCub with existing iCub simulators (iCub SIM [26], iCub Gazebo [17], and neuromorphic iCub [6]). *Skin to Gazebo was added later by the community. **The planners are coming from the connection with YARP. ***By default, can be changed.****Event-driven skin.
@@
Middleware
Transferable
to real iCub
Skin
High-level
planners
Physics
Engine
Rendering
Engine
iCub SIM
@@
YARP Yes Yes* Yes** DART*** OGRE
Neuromorphic
iCub
Python None No Yes**** No MuJoCo OpenGL
pyCub Python None No Yes No Bullet
@@
We present pyCub, a Python physics-based simulator of the humanoid robot iCub. The simulator contains all limbs of the robot, together with the possibility to visualize view from the robot’s cameras, and with simulated artificial whole-body skin consisting of more than 4000 individual tactile sensors. The framework provides easy-to-use functions to control the robot in joint (position and velocity) and Cartesian space, and effortless access to proprioceptive, visual, and tactile sensors of the robot. Everything is written and accessible in Python, allowing even less advanced users to operate. The environment is fully customizable and allows to insert other objects—see Fig. 1 for an example. In addition to the simulator, we provide exercises of different difficulty levels ranging from basic control and movements to tasks typical for humanoid robots such as gazing, grasping, or whole-body skin processing and obstacle avoidance. The exercises can be further customized and allow difficulty adjustments and thus are suitable for learning (humanoid) robotics on different difficulty levels. The framework was deployed in two runs of a course on humanoid robotics.
@@ 2 Simulation Environment
2.1 Technical Details
The physics-backbone of the simulator is realized with the PyBullet engine [4]. This module was selected because it is open-source, multi-platform and does not require excessive computational resources—mainly compared to recent physical simulators such as Orbit (NVIDIA Issac Sim) [18]. The schematic of the simulator package is shown in Fig. 2.
We tested the performance (the ratio of simulation time and real time; the bigger the better) on an average laptop and a workstation, and the results are in Table 2. We tested two models: (i) a full model with detailed meshed and full fingers; (ii) a simplified model with less detailed meshes and without fingers. We can see that on a laptop for the full model with skin and Graphical User Interface (GUI) enabled, the simulation runs around real-time (0.95 in this case, but the performance is based on the number of skin events). With the simple model, it increases to around 1.5. This performance is enough for the use case of the framework as one wants to see the robot moving in real-time or even slower. Without skin, performance is always higher than real time. For the workstation, the performance is always higher—PyBullet uses a single core for all computations, and the efficiency of single-core computations makes the difference between the laptop and the workstation. In addition, tasks such as reinforcement learning could be achieved by running more instances on more cores.
Table 2: Average real-time factors (the ratio of simulation time and real time; the bigger the better) for two models (full model (FM) and a simple model (SM)) for four configurations—with and without the GUI with and without the skin enabled. The results are averaged over 10 repetitions. Tested on an average laptop CPU (LAP) Ryzen 7 PRO 4750u and a workstation CPU (WS) Intel i9-11900.
GUI Headless GUI with skin Headless with skin
FM - LAP $1.37\pm 0.02$ $2.03\pm 0.02$ $0.95\pm 0.01$ $1.07\pm 0.02$
SM - LAP $6.73\pm 0.07$ $8.95\pm 0.11$ $1.46\pm 0.05$ $1.90\pm 0.08$
FM -WS $2.25\pm 0.00$ $2.41\pm 0.00$ $1.29\pm 0.01$ $1.41\pm 0.00$
SM - WS $0.82\pm 0.27 11$$1.46\pm 0.03$ $2.58\pm 0.02$ $2.77\pm 0.02$
2.1.1 Time Management
The simulation is not running by itself, but the users have to perform every step manually, i.e., call update_simulation(). It is advantageous because even when using less powerful computers, there is enough time to check the state of the world and perform computations between simulation steps, i.e., the visualization may run slower but everything will be computed correctly. Additionally, there is an option to actually perform the simulation step only if a predefined duration elapsed from the last one. It is useful to slow down the time to, for example, check movements in more detail. The GUI has its own fixed time management, as it is required to catch mouse events—zoom, rotations, etc.
2.1.2 GUI
@@
Figure 3: Example of native GUI with output from both eyes and additional visualization showing vectors where the robot is looking (blue) and where it should be looking (red) used in the gaze exercise.
2.1.3 Customizations
The environment is fully customizable through .yaml configuration files. The files are loaded when initializing an instance of the simulator, and thus users can create different configurations for different exercises/tasks/demos. The basic options allow to enable/disable GUI, artificial skin, eye-cameras output, logging, or self collisions. Further, the configurations allow us to set an initial position for every joint of the robot, the end-effector used for Cartesian control, or simulation step size. Finally, users can insert additional objects into the simulation. The configurations support Unified Robot Description Format (URDF) files, or directly .obj triangular-mesh files. One can further specify the initial position, color, and whether the object is subject to gravity and can be moved (e.g, ball) or not (e.g., table).
2.1.4 Installation
To provide easy installation, the framework can be installed from PyPI. Alternatively, we provide several versions of prebuilt Docker images or a version runnable entirely online using the GitPod [^1] https://www.gitpod.io/ platform. Everything is available at https://rustlluk.github.io/pyCub.
2.2 Control Modes
@@
Joint Velocity Control. The most direct and low-level is the velocity control. It directly sets the joint velocity (in  $\text{\,}\mathrm{rad}\text{\,}{\mathrm{s}}^{-1}$) of every controlled joint. It is useful for reactive control. In this mode, the robot moves until a different command is set by the user.
Joint Position Control. The position control sets the target joint position of each controlled joint (in  $\text{\,}\mathrm{rad}$). In this mode, the robot moves until all joints reach their set-point or, optionally, until a collision is detected. The users can additionally select whether to care about self-collision—collisions of parts of the robots.
@@
2.3 Artificial Skin
The artificial skin is an integral component of the robot. However, the real robot contains more than 4000 individual sensors, making the simulation challenging. The sensors are also not binary, i.e., they can return value with 8-bit resolution. We decided to simulate the skin using raycasting. From every skin sensor (taxel) we cast a short ray with a given maximum length. If the ray hits something (other part of the robot or object in the environment), the output of the sensor is computed proportionally to the length of the ray. Using this method, we are able to simulate the intensity of contact and also the fact that for the real skin, even taxels around the touch point are lightly activated because of the deformation of the skin. The raycasting needs to be done in every step of the simulation, and, even though the raycasting is optimized and done in batch, it is still computationally demanding. Thus, we decided to first check (for every skin part) whether there is a close object (using overlaps of bounding boxes of objects) and perform the raycasting only if necessary. The values for each activated taxels are saved in a dictionary inside the simulator and binary activations are shown as green points in the visualization—see Fig. 4.
@@ 2.3 Artificial Skin
3 Exercises
This Section introduces and describes the exercises available in the framework. We prepared several exercises that, in our opinion, span important tasks in robotics. We will describe those in recommended order based on their difficulty. However, the exercises can be further adjusted to change the difficulty level. The exercises always contain a configuration file, an assignment script, and a testing script. Videos of each exercise are available at https://rustlluk.github.io/pyCub and in the accompanying video.
3.1 Push the Ball!
The first exercise is purely for the purpose of learning how to control the simulator. The goal is to get a ball as far from a table as possible—see Fig. 1 for visualization. Basically, every imaginable approach is allowed. Whether hitting, pushing, or grasping and throwing the ball. Upon completion, users should be able to control the robot in some of the modes and understand basic functionalities. During this exercise, the self-collisions are turned off—it is visible in some examples in the accompanying video.
The testing script for this exercise computes the distance of a ball and a table after two seconds after the last movement of the robot and provides the distance in meters.
3.2 Smooth Movements
As described in Section 2.2, the Cartesian interface of the simulator does not include any underlying planner. This exercise aims to provide insight on how to move in a straight line or how to “draw” a shape, e.g., a circle. In addition, the movements should be “smooth”, i.e., a continuous movement without any jerks.
@@ 3.2 Smooth Movements
The testing script computes the distance from the trajectory performed to a reference one. There are different metrics for line and circle movements. However, the smoothness of the movements must be checked manually for now.
3.3 Gaze
Gaze control is a key ability so that the robot can bring some objects into the field of view. The goal of gaze can be described as focusing on the object of interest using, mainly, joints in the head and eyes. There is a gaze control implemented in the real iCub robot [22] that is capable of fixing to a given 3D using all joints in the neck and eyes. In this exercise, the problem is simplified to fixating the view of the robot using neck joints to a ball moving in XY-plane.
@@ 3.3 Gaze
The described solution is only one of many and can be solved in both simpler and more complicated ways. In addition, the thresholds for the allowed error can be adjusted, which can make some simpler solutions infeasible.
@@
3.4 Reactive Control and Skin Processing
The fourth exercise focuses on reactive control in the case of contact with the environment. It is an important aspect of Human-Robot Interaction (HRI), since ideal humanoid robots should be able to operate in close contact with humans, which inevitably leads to collisions.
@@ 3.4 Reactive Control and Skin Processing
where $\mathbf{J(q)}^{*}$ can be either a transposition $\mathbf{J(q)}^{T}$, an inverse $\mathbf{J(q)}^{-1}$, or a Moore-Penrose pseudoinverse $\mathbf{J(q)}^{+}$. One of the educational points of the exercise is to test the options, as each of them has its pros and cons. The classic inverse can only be used when the Jacobian is square and has full rank. Pseudoinverse can be unstable near singular configurations and move the end-effector in a straight line. Transpose is easier to compute, but does not necessarily move in a straight line. In the exercise, some limbs are in a configuration in which not all options provide the desired movement.
The exercise has different levels of difficulty depending on the limb in contact and the number of simultaneous contacts. The testing script here runs the different levels of difficulty. At the moment, there is no automatic evaluation and performance must be checked manually.
3.5 Grasping
Another important skill for humanoids is object handover between the human and the robot. The robot must be reactive (the previous exercise) and must be able to grasp objects. Robotic grasping has changed a lot during its development. From grasping based on friction cone and wrenches, through physics-based grasping to Machine Learning (ML) and Deep Learning (DL)-based approaches [5].
This exercise shows “core” grasping and is simplified by the fact that we always want to grasp the same green ball—the only thing that changes is the position of the ball on a table. The aim is also to practice core computer vision, and thus the users are guided to find the ball in the image plane (in an RGB image from one of the eyes) and obtain 2D center of the ball. The center can be deprojected to 3D given the center (u, v pixels in an image), the depth value for the tuple of pixels, and camera parameters (width; height; focal length). Then, the robot should move above the ball, get fingers to correct start position, move toward the ball, grasp it, and move it up to prove a stable grasp. An example of successful grasp can be seen in Fig. 1.